22) There is an order for 1 Liter D5 ½ NS to infuse over 12 hrs a) How many mls total per day is needed? b) Available is a maPlease answer parts a-c. Thank you!
Answer #1

ORDER IS

1 LITRE D5 AND HALF NS

1 LITRE D5 = 1000 ML

HALF NS ( 0.45 %) = 1000 ML

TOTAL VOLUME IS 2000 ML

     TIME = 12 HOURS

ANSWER ( A)

1 DAY = 24 HOURS

FOR 12 HOURS = 2000ML ( GIVEN)

SO, FOR 24 HOURS = 2 *2000 ML

= 4000ML    PER 24 H OURS

     TOTAL 4000 ML IS NEEDED FOR 24 HOURS

ANSWER (B)

MICRODROP IS USED TO ADMINISTEr SMALLER DOSE AND CERATIN MEDICATIONS SUCH AS ANTIBIOTICS, OR OTHER DRUGS.

MACRODRIP IS USED TO ADMINISTER ROUTINE FLUIDS FOR LONGER PERIOD OF TIME

SO, IN THIS CONDITION MACRODRIP WILL BE USED, AS THERE IS A NEED TO ADMINISTER FLUIDS FOR A LONGER PERIOD OF TIME, AND ALSO MICRODRIP IS USED TO ADMINISTER FLUIDS IN PEDIATRIC PATIENT AND MACRO DRIP IS USED IN ADULT.

ALSO, THESE SOLUTIONS ARE HYPOTONIC AND CAN BE USED VIA MACRODRIP.

ANSWER (C)

FLOW RATE (DROPS/MIN) = TOTAL VOLUME OF FLUID (ml) * DROP FACTOR / TOTAL TIME OF INFUSION (minutes)

= 2000 (ml) * 10 / 12 *60

= 20000 / 720 (min)

= 27.77777 DROPS / MIN

= 28 DROPS/ MIN WHEN MACRODRIP IS USED.